1. In large department store there are 500 employees. 350 are famel and 200 them are under the age of 25. There are 75 males under 25 if an employee’s selected for promotion find the probability that the employees will be the following.
a. Under 25 or a female
b. Over 24 or female
c. Male over 24

Answer by Boreal(15235)

Gender----M -----F-----Total
Age <25---75----125----200
Age >24---75 ---225---300
----------150----350---500
under 25 OR a female is 200 under 25+350 female-both (125)=550-125=425,probability is 425/500=0.85
over 24 or a female is 300 over 25+350 female minus 225 both which is 425/500 with probability 0.85
Male over 24 is 75/500 with probability of 0.15.